# Unsourced Adversarial CAPTCHA: A Bi-Phase Adversarial CAPTCHA Framework

### v2

@@ IV METHODOLOGY
In this section, we introduce the UAC framework in the context of white-box attacks, where we integrate both the LLM and EDICT frameworks to generate unsourced adversarial examples while ensuring the quality of the generated images. Subsequently, to address the broader applicability and higher complexity of black-box attack scenarios, we further extend the UAC framework by incorporating the clip model and the bi-path optimization strategy. This enhancement ensures semantic consistency during the diffusion model’s generation process and overcomes the limitations of traditional search spaces, enabling the generation of unsourced adversarial CAPTCHAs suitable for practical application scenarios. The overall framework is shown in Fig. 2.
@@ IV-A Threat model
Considering the distinct contexts of image recognition and CAPTCHA security, we adopt a threat model based on image classification and examine both white-box and black-box attack scenarios. In this scenario, the attacker only needs to provide the target class and the desired adversarial instance class during the generation phase and manipulate the test image during the inference phase without involvement in the training phase of the model.
IV-B Unsourced adversarial CAPTCHA
@@ IV-B1 Prompt guidance
Firstly, we introduce an LLM to ensure the accuracy and diversity of the generation process. Although the initial attack aims to generate the corresponding adversarial examples based on the object categories entered by the attacker, in the experiments, we found that the use of only short prompts $P^{\prime}$ tends to lead to generation errors. This is mainly because brief prompts lack sufficient contextual information, making the diffusion model ambiguous in understanding the target image features. Diffusion models rely on the semantics of the prompter to gradually guide image generation. Still, when the prompter is too simple, for example, a single word or an ambiguous description, the model may associate it with multiple different potential image features, resulting in generation results that deviate from expectations.
To address this issue, we introduce an LLM [21, 22, 23] to enhance the semantic information of the prompt. The LLM guides the generator’s inputs by generating richer and contextually relevant prompts like “A goldfish swims in the bowl”, which not only ensures the diversity and stability of the generated examples but also enables the attacker to generate more accurate adversarial examples given the category and the specific target. That is :
$P^{\prime}=f_{\mathrm{LLM}}(P)$ (8)
Next, we use a diffusion model $G$ to generate images step by step, and the generation process of this model is conditionally guided by the extended prompt. Specifically, we first input the extended prompt $P^{\prime}$ into $G$ for conditional generation, and $G$ generates a representation of the intermediate latent variable $z_{t}$ based on the latent variables and the prompt at each time step $t$. Specifically, this process can be formulated as
$I_{t}=G(z_{t},P^{\prime},\epsilon_{t})$ (9)
where $\epsilon_{t}$ is the noise term we introduce in the generation process $t$ to ensure diversity in the generation.
Algorithm 1 Adversarial example synthesis in UAC
@@
12:        $y^{inter}_{t-1}=y_{t}-\eta_{y}*\nabla_{y_{t}}\ell(y_{t},y_{target})$
13:        Merge optimized latent variables into a unified representation:$z_{t-1}=\alpha*x^{inter}_{t-1}+(1-\alpha)*y^{inter}_{t-1}$
14:        $I_{t-1}=G(z_{t-1})$
@@ IV-B2 Latent pairwise diffusion optimization
In our approach, gradient guidance and merging of latent variables are key steps to ensure that the representation of latent spaces during generation accurately guides the optimization of the generator $G$. Specifically, the gradient of the target model $f$ needs to be used first to guide the latent variable pairs:
$x_{t-1}^{\mathrm{inter}}=x_{t}-\eta_{x}\nabla_{x_{t}}\ell(x_{t},y_{target})$ (10)
$y_{t-1}^{\mathrm{inter}}=y_{t}-\eta_{y}\nabla_{y_{t}}\ell(y_{t},y_{target})$
where $L$ denotes the loss function, $\nabla$ indicates the gradient of $f$, $y_{target}$ indicates the attacker specifies the target label of the attack, and $\eta$ denote the learning rate of the latent variable pairs $(x_{t},y_{t})$
@@
$\begin{gathered}z_{t-1}=\alpha\cdot x_{t-1}^{\mathrm{inter}}+(1-\alpha)\cdot y_{t-1}^{\mathrm{inter}}\\ \begin{aligned} \mathrm{s.t.}\>\alpha\in[0,1]\end{aligned}\end{gathered}$ (13)
where $\alpha$ is a weight parameter to control the proportion of the contribution of the two latent variables in the merger. The goal of the merging process is to fuse the optimized properties of the two latent spaces to produce a more robust and integrated latent representation that better supports the generation of diffusion models.
@@ IV-B3 Adversarial example synthesis
In UAC, $P^{\prime}$ is provided to the generator G as the input and initial condition for the adversarial example generation process. During the reverse diffusion process of the generator, EDICT is employed as a reversible temporal diffusion framework to iteratively optimize the latent variable pair $(x_{t},y_{t})$. At each step, gradient guidance is utilized to progressively refine the generated image toward the target class $y_{target}$. Specifically, for each latent variable $z_{t}$ in the generation process, joint optimization of $(x_{t},y_{t})$ is performed, leveraging the gradient information of the target loss to guide the latent variable closer to the target class during optimization. At each time step $t$ UAC adjusts the variable pair $(x_{t},y_{t})$ through gradient guidance, controlling the diffusion and denoising processes. Ultimately, the synergy between gradient guidance and latent variable optimization ensures that the generated adversarial examples not only precisely align with the target conditions but also achieve optimal performance in terms of distribution consistency and attack efficacy. In summary, our adversarial example generation can be expressed by the following equation:
$z^{\prime}=\arg\min_{z_{t}}\ell(z_{t},P,y_{target})$ (14)
@@
where $z^{\prime}$ denotes the final latent variant. After this, we pass the final latent variable $z^{\prime}$ to the generator $G$ to obtain the final output of the adversarial examples $I^{*}=G(z^{\prime})$. The specific adversarial example synthesis process is shown in algorithm 1.
IV-C Bi-path unsourced adversarial CAPTCHA
@@
Although UAC can effectively deceive the machine into producing false recognition results by utilizing the original model information in a white-box environment, the premise of a white-box environment is that the attacker has access to the complete structure and parameters of the target model. However, in practical applications, especially in adversarial CAPTCHA scenarios, this premise often does not hold because the detailed information of the target model is not available, and direct access to the original image source is also challenging due to the copyright issues of the image source. Therefore, we extend the UAC method into a more practical black-box attack method called BP-UAC.
In BP-UAC, before the attack process, we introduce the gradient of the clip model [24] to guide $G$ to generate the initial states $x_{m}$ and $y_{m}$, and synthesize the two initial states obtained into a temporary latent variable $z_{m}$ through Eq. 13 after the conclusion of the bootstrap optimization of the clip model. To ensure the consistency of the generated image with the $P^{\prime}$.
This consistency not only enhances the naturalness and visual quality of generated images but also prevents distribution drift. Additionally, incorporating clip model’s gradient significantly improves the stealthiness of adversarial examples by aligning the generated images with the prompt description, making them harder for human observers and detection methods to identify.
@@
Then, we integrate the gradients of multiple models during the attack and propose a bi-path optimization strategy to address the limitations of white-box environments and adapt to black-box scenarios. By merging predictions from multiple models, we can approximate the behavior of the target model, enhancing the ASR on unknown models. The bi-path strategy balances the loss from the target class and second-highest posterior probabilities in the target model’s output, effectively capturing vulnerabilities in the decision boundary to improve adversarial robustness and ASR.
@@
Compared to traditional attacks, BP-UAC excels in generating perturbations by exploring the input space more comprehensively through bi-path optimization strategy. This increases the diversity of adversarial examples and enhances the attack’s robustness, allowing it to exploit weaknesses in the target model more effectively. As a result, BP-UAC achieves higher success rates in complex and uncertain black-box environments.
@@
To accomplish the above, we need to add the assumption that we have three known proxy models $f_{\mathrm{1}}$ ,$f_{\mathrm{2}}$ and $f_{\mathrm{3}}$ , which have classification predicted probability distribution of $g_{1}(x)$, $g_{2}(x)$, $g_{3}(x)$, and parameters $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$, respectively. Using the predict probability distribution of these models, we update the latent variables $x_{t}$ and $y_{t}$. Based on UAC, we replace the known model gradients in Eq. 10 and Eq. 13 with the weighted average of the gradients from the three proxy models to enhance the attack’s effectiveness against an unknown black-box model. Specifically, our integrated classification predicted probability distribution $g_{x_{t}}$ and $g_{y_{t}}$ can be expressed by the following equation:
@@
$g_{x_{t}}=\frac{\beta_{1}g_{1}(x)+\beta_{2}g_{2}(x)+\beta_{3g_{3}(x})}{\beta_{1}+\beta_{2}+\beta_{3}}$
where $g_{1},g_{2},g_{3}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{M}$, the calculation of $g_{y_{t}}$ is the same as for $g_{x_{t}}$.
Algorithm 2 Adversarial example synthesis in BP-UAC
@@
3:  Using Eq. 9 and Eq. 10 optimize the latent variables $x$ and $y$ through $m$, get $x_{m}$ and $y_{m}$
@@
12:        Integrate $\ell_{div}$, $\ell_{tar}$ as $\ell_{x_{t}}$ base on Eq. 14
13:        Compute gradient $\nabla$ base on Eq. 15
@@
15:        $x^{inter}_{t_{1}}=x-\eta_{x}*\nabla_{x_{t}}\ell_{x_{t}}$
16:        $y^{inter}_{t-1}=y-\eta_{y}*\nabla_{y_{t}}\ell_{y_{t}}$
17:        Merge optimized latent variables into a unified representation:$z_{t-1}=\alpha*x^{inter}_{t-1}+(1-\alpha)*y^{inter}_{t-1}$
18:        $I_{t}=G(z_{t-1})$
@@
Leveraging the collective gradient information from these proxy models, we further guide the model to generate adversarial examples toward the second-highest and target class probability classes. The bi-path optimization strategy identifies a more robust path across different loss spaces, enhancing the transferability and success rate of adversarial examples in black-box models. This approach effectively improves the performance of black-box attacks, making adversarial examples more deceptive and robust.
Specifically, we focus on guiding the generation process by minimizing the loss associated with both the second-highest and target class probability classes:
$\ell_{div}$ $=-\ell(x_{t},y_{second})$ (16)
$\ell_{tar}$ $=\ell(x_{t},y_{target})$
Finally, the new loss $\ell_{BP}$ is derived by aggregating the losses from multiple directional objectives, effectively balancing the guidance effects of both the target and auxiliary classes during optimization, thereby overcoming the challenges associated with black-box model attacks:
$\ell_{x_{t}}$ $=\frac{\ell_{div}+\ell_{tar}}{2}$ (17)
@@
$\nabla_{x_{t}}=\frac{\partial\ell_{x_{t}}}{\partial g{x_{t}}}$ (18)
@@
Surrogate Model: RN-50 Method
Target Model SAE ADer ReColorADV cAdv tADV NCF ACE ColorFool ACA Ours
@@
Average 49.05 17.75 39.25 45.9 41.7 30.9 32.75 56.3 67.2 81.9
Surrogate Model: MN-v2 Method
Target Model SAE ADer ReColorADV cAdv tAdv NCF ACE ColorFool ACA Ours
@@
BIM TSR 75.3% 67.9% 62.3% 53.6% 81.7% 83.5% 64.3%
UAC USR 100% 100% 100% 100% 100% 100% 100%
UAC TSR 100% 100% 100% 100% 99% 100% 99%
TABLE V: The performance comparison of BP-UAC and some state-of-art adversarial attack methods relying on generative models on ASR.
Baseline Models Attack method : BP-UAC
@@ V Experiments and Results
In this section, we will first validate the attack effectiveness of our method in white-box and black-box scenarios, respectively, by comparing the attack effectiveness with other methods and calculating the attack effectiveness of existing methods in the face of unknown models. Next, we launch a comprehensive ablation experiment to compare the attack effect, the quality of the generated images, and the efficiency of the attack, respectively, to prove the usefulness of our module and the rationality of the parameter settings.

### v3

@@ IV METHODOLOGY
In this section, we introduce the DAC framework in the context of white-box attacks, where we integrate both the LLM and EDICT frameworks to generate unsourced adversarial examples while ensuring the quality of the generated images. To address the broader applicability and higher complexity of black-box attack scenarios, we further extend the DAC framework by incorporating the clip model and the bi-path optimization strategy. This enhancement ensures semantic consistency during the diffusion model’s generation process and overcomes the limitations of traditional search spaces, enabling the generation of unsourced adversarial CAPTCHAs suitable for practical application scenarios. The overall framework is shown in Fig. 2.
@@ IV-A Threat model
Considering the distinct contexts of image recognition and CAPTCHA security, we adopt a threat model based on image classification and examine both white-box and black-box attack scenarios. In this scenario, the attacker only needs to provide the target class and the desired adversarial instance class during the generation phase, and manipulate the test image during the inference phase without involvement in the training phase of the model.
IV-B Defensive adversarial CAPTCHA
@@ IV-B1 Prompt guidance
First, we introduce an LLM to ensure the accuracy and diversity of the generation process. Although the initial attack aims to generate corresponding adversarial examples based on the object categories entered by the attacker, our experiments reveal that using only short prompts $P^{\prime}$ tends to lead to generation errors. This is because brief prompts lack sufficient contextual information, making the diffusion model ambiguous in understanding the target image features. Diffusion models rely on the semantics of the prompter to gradually guide image generation. When the prompt is too simple, for example, a single word or an ambiguous description, the model may associate it with multiple different potential image features, resulting in generation results that deviate from expectations.
To address this issue, we introduce an LLM [37, 38, 39] to enhance the semantic information of the prompt. The LLM guides the generator’s inputs by generating richer and contextually relevant prompts like “A goldfish swims in the bowl”, which not only ensures the diversity and stability of the generated examples but also enables the attacker to generate more accurate adversarial examples, given the category and the specific target. That can be written as:
$P^{\prime}=f_{\mathrm{LLM}}(P).$ (8)
Next, we utilize a diffusion model $G$ to generate images step by step, with the generation process of this model conditionally guided by the extended prompt. Specifically, we first input the extended prompt $P^{\prime}$ into $G$ for conditional generation, and $G$ generates a representation of the intermediate latent variable $z_{t}$ based on the latent variables and the prompt at each time step $t$. Specifically, this process can be formulated as:
$I_{t}=G(z_{t},P^{\prime},\epsilon_{t}),$ (9)
where $\epsilon_{t}$ is the noise term that we introduce in the generation process $t$ to ensure diversity in the generation.
Algorithm 1 Adversarial example synthesis in DAC
@@
12:        $y^{inter}_{t-1}=y_{t}-\eta_{y}*\nabla_{y_{t}}\ell(y_{t},y_{target})$;
13:        Merge optimized latent variables into a unified representation:$z_{t-1}=\alpha*x^{inter}_{t-1}+(1-\alpha)*y^{inter}_{t-1}$;
14:        $I_{t-1}=G(z_{t-1})$;
@@ IV-B2 Latent pairwise diffusion optimization
Gradient guidance and merging of latent variables are key steps to ensure that the representation of latent spaces during generation accurately guides the optimization of the generator $G$. Specifically, the gradient of the target model $f$ needs to be used first to guide the latent variable pairs:
$x_{t-1}^{\mathrm{inter}}=x_{t}-\eta_{x}\nabla_{x_{t}}\ell(x_{t},y_{target});$ (10)
$y_{t-1}^{\mathrm{inter}}=y_{t}-\eta_{y}\nabla_{y_{t}}\ell(y_{t},y_{target}),$
where $L$ denotes the loss function, $\nabla$ indicates the gradient of $f$, $y_{target}$ indicates the attacker specifies the target label of the attack, and $\eta$ denotes the learning rate of the latent variable pairs $(x_{t},y_{t})$.
@@
$\begin{gathered}z_{t-1}=\alpha\cdot x_{t-1}^{\mathrm{inter}}+(1-\alpha)\cdot y_{t-1}^{\mathrm{inter}},\end{gathered}$ (12)
where $\alpha\in[0,1]$ is a weight parameter to control the ratio of the contribution of the two latent variables in the merger. The goal of the merging process is to combine the optimized properties of the two latent spaces, producing a more robust and integrated latent representation that better supports the generation of diffusion models.
@@ IV-B3 Adversarial example synthesis
In DAC, $P^{\prime}$ is provided to the generator G as the input and initial condition for the adversarial example generation process. During the reverse diffusion process of the generator, we leverage EDICT to design parameter-shared coupled noise variable pairs $(x_{t},y_{t})$ in the diffusion model generation process and achieved robust inversion through an iterative bidirectional calibration mechanism at discrete time steps. In each diffusion step, one of the variables calculates the denoising direction based on the state of the other variable, and the result of this calculation is immediately incorporated into the conditional update input of the other vector. This cross-feedback mechanism constructs a dynamic error correction system. This method approximates the initial input by performing one denoising operation at each time step, enabling the model gradient to propagate backward through the entire chain. At each step, gradient guidance is utilized to progressively refine the generated image toward the target class $y_{target}$. Specifically, for each latent variable $z_{t}$ in the generation process, joint optimization of $(x_{t},y_{t})$ is performed, leveraging the gradient information of the target loss to guide the latent variable closer to the target class during optimization. At each time step $t$, DAC adjusts the variable pair $(x_{t},y_{t})$ through gradient guidance, controlling the diffusion and denoising processes. The synergy between gradient guidance and latent variable optimization ensures that the generated adversarial examples not only precisely align with the target conditions, but also achieve optimal performance in terms of distribution consistency and attack efficacy. Our adversarial example generation can be mathematically expressed as:
$z^{\prime}=\arg\min_{z_{t}}\ell(z_{t},P,y_{target}),$ (13)
@@
where $z^{\prime}$ denotes the final latent variable. After this, we pass $z^{\prime}$ to the generator $G$ to obtain the final output of the adversarial examples $I^{*}=G(z^{\prime})$. The specific adversarial example synthesis process is summarized in Algorithm 1.
IV-C Bi-path optimization strategy
@@
Although DAC can effectively deceive the machine into producing false recognition results by utilizing the original model information in a white-box environment, the premise of a white-box environment is that the attacker has access to the complete structure and parameters of the target model. In practical applications, especially in adversarial CAPTCHA scenarios, this premise often does not hold because the detailed information of the target model is not available, and direct access to the original image source may not be available due to the copyright of the image source. Therefore, we extend the DAC method into a more practical black-box attack method, called BP-DAC.
In BP-DAC, prior to the attack phase, we leverage the CLIP model’s gradient [40] to guide generator G in producing the initial states $x_{m}$ and $y_{m}$; following CLIP’s bootstrap optimization, these states are synthesized via Eq.(12) into a temporary latent variable $z_{m}$. This CLIP-guided initialization critically enforces consistency between the generated image and the target distribution $P^{\prime}$, enhancing naturalness, visual quality, and preventing distribution drift, while significantly bolstering adversarial stealth through alignment with the prompt description, thus evading both human and automated detection. During the attack phase, we integrate gradients from multiple models and introduce a bi-path optimization strategy. This strategy addresses white-box limitations and enhances black-box adaptability by approximating the target model’s behavior through merged predictions, thereby boosting the Attack Success Rate (ASR) on unknown models. The bi-path component specifically balances losses derived from the target class and the second-highest posterior probabilities, effectively exploiting vulnerabilities at the decision boundary to jointly enhance adversarial robustness and ASR.
@@
Then, we integrate the gradients of multiple models during the attack and propose a bi-path optimization strategy to address the limitations of white-box environments and adapt to black-box scenarios. By merging predictions from multiple models, we can approximate the behavior of the target model, enhancing the ASR on unknown models. The bi-path strategy balances the loss from the target class and the second-highest posterior probabilities in the target model’s output, effectively capturing vulnerabilities in the decision boundary to improve adversarial robustness and ASR.
@@
Compared to traditional attacks, BP-DAC excels in generating perturbations by exploring the input space more comprehensively through a bi-path optimization strategy. This increases the diversity of adversarial examples and enhances the attack’s robustness, allowing it to effectively exploit weaknesses in the target model. As a result, BP-DAC achieves higher success rates in complex and uncertain black-box environments.
@@
To accomplish the above, we need to add the assumption that we have three known proxy models $f_{\mathrm{1}}$,$f_{\mathrm{2}}$ and $f_{\mathrm{3}}$, which have classification predicted probability distribution of $g_{1}(x)$, $g_{2}(x)$, $g_{3}(x)$, and the parameters $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$, respectively. Using the predict probability distribution of these models, we update the latent variables $x_{t}$ and $y_{t}$. Based on DAC, we replace the known model gradients in Eq.(10) and Eq.(12) with the weighted average of the gradients from the three proxy models to enhance the attack’s effectiveness against an unknown black-box model. Specifically, our integrated classification predicted probability distribution $g_{x_{t}}$ and $g_{y_{t}}$ can be expressed as:
@@
$g_{x_{t}}=\frac{\beta_{1}g_{1}(x)+\beta_{2}g_{2}(x)+\beta_{3g_{3}(x})}{\beta_{1}+\beta_{2}+\beta_{3}},$
where $g_{1},g_{2},g_{3}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{M}$, and the calculation of $g_{y_{t}}$ is the same as for $g_{x_{t}}$.
Algorithm 2 Adversarial example synthesis in BP-DAC
@@
3:  Using Eq. (9) and Eq. (10) optimize the latent variables $x$ and $y$ through $m$, get $x_{m}$ and $y_{m}$
@@
12:        Integrate $\ell_{div}$, $\ell_{tar}$ as $\ell_{x_{t}}$ base on Eq. (14)
13:        Compute gradient $\nabla$ base on Eq. (15)
@@
15:        $x^{inter}_{t_{1}}=x-\eta_{x}*\nabla_{x_{t}}\ell_{x_{t}}$;
16:        $y^{inter}_{t-1}=y-\eta_{y}*\nabla_{y_{t}}\ell_{y_{t}}$;
17:        Merge optimized latent variables into a unified representation:$z_{t-1}=\alpha*x^{inter}_{t-1}+(1-\alpha)*y^{inter}_{t-1}$;
18:        $I_{t}=G(z_{t-1})$;
@@
Leveraging the collective gradient information from these proxy models, we further guide the model to generate adversarial examples toward the second-highest and target class probabilities. The bi-path optimization strategy identifies a more robust path across different loss spaces, enhancing the transferability and success rate of adversarial examples in black-box models. This approach effectively improves the performance of black-box attacks, making adversarial examples more deceptive and robust.
Specifically, we focus on guiding the generation process by minimizing the loss associated with both the second-highest and target class probabilities:
$\ell_{div}$ $=-\ell(x_{t},y_{second});$ (15)
$\ell_{tar}$ $=\ell(x_{t},y_{target}).$
Finally, the new loss $\ell_{BP}$ is derived by aggregating the losses from multiple directional objectives, effectively balancing the guidance effects of both the target and auxiliary classes during optimization. This overcomes the challenges associated with black-box model attacks, as given by:
$\ell_{x_{t}}$ $=\frac{\ell_{div}+\ell_{tar}}{2}.$ (16)
@@
$\nabla_{x_{t}}=\frac{\partial\ell_{x_{t}}}{\partial g{x_{t}}}.$ (17)
@@
Target Method
Model SAE ADer ReColorADV cAdv tADV NCF ACE ColorFool ACA DAC
@@
[…]
Target Method
Model SAE ADer ReColorADV cAdv tAdv NCF ACE ColorFool ACA DAC
@@
TSR 75.3% 67.9% 62.3% 53.6% 81.7% 83.5% 64.3%
DAC USR 100% 100% 100% 100% 100% 100% 100%
TSR 100% 100% 100% 100% 99% 100% 99%
TABLE V: The performance comparison of BP-DAC and some state-of-art adversarial attack methods relying on generative models on ASR.
Baseline Models Attack method : BP-DAC
@@ V Experiments and Results
In this section, we first validate the effectiveness of our method in white-box and black-box scenarios by comparing it with other methods and calculating the attack effectiveness of existing methods in the face of unknown models. Next, we conduct a comprehensive ablation experiment to compare the attack effect, the quality of the generated images, and the efficiency of the attack, thereby proving the usefulness of our module and the rationality of our parameter settings.